# RTop-K: Ultra-Fast Row-Wise Top-K Selection for Neural Network Acceleration on GPUs

### v3

@@ 2.3 GPU Top-k Implementations
4:  $thres,cnt\leftarrow max,0$
5:  while $thres-min>\epsilon$ do
6:     $cnt\leftarrow\left|\{i\mid\mathbf{v}_{i}\geq thres\}\right|$
7:     if $cnt<k$ then
8:        $max\leftarrow thres$
9:     else if $cnt>k$ then
10:        $min\leftarrow thres$
11:     else
12:        break
13:     end if
14:     $thres\leftarrow\frac{min+max}{2}$
@@
16:  $borderline\_cnt\leftarrow k-cnt$
17:  if $borderline\_cnt\geq 0$ then
18:     $\mathbf{elems},\mathbf{indices}\leftarrow\{(\mathbf{v}_{i},i)\mid\mathbf{v}_{i}\geq thres\}$
19:     Append the first $borderline\_cnt$ pairs of $\{(\mathbf{v}_{i},i)\mid min\leq\mathbf{v}_{i}<thres\}$ to $\mathbf{elems},\mathbf{indices}$
20:  else
21:     $\mathbf{elems},\mathbf{indices}\leftarrow$ the first $k$ pairs of $\{(\mathbf{v}_{i},i)\mid\mathbf{v}_{i}\geq thres\}$
22:  end if
23:  return  $\mathbf{elems},\mathbf{indices}$
@@ 3.1 Binary Search-based Top-k Selection Algorithm
The algorithm, as illustrated in Fig. 2, first retrieves the $min$ and $max$ values of the vector, and then uses several iterations of binary search to determine a threshold $thres$. The $min$, $max$, and $thres$ values are updated in each iteration, and the loop terminates when the number of elements filtered by the current threshold equals $k$. One corner case that must be considered is when $\mathbf{v}$ contains multiple equal or very close elements, and these elements happen to be near the ”borderline” during top-k selection. In this scenario, relying solely on $thres$ for selection makes it difficult or even impossible to extract exactly $k$ elements. It can be verified that in such cases, these borderline elements are always located between $min$ and $thres$. Therefore, a second filtering step using $min$ is sufficient to supplement the selection and ensure that exactly $k$ elements are chosen. The parameter $\epsilon$ represents the precision of the algorithm, which determines the maximum width of the borderline and imposes an upper bound on the number of iterations in the loop. When $\epsilon$ is set to 0, Algorithm 1 becomes an exact top-k selection algorithm.
@@ 3.2 GPU Implementation Design
Both Algorithm 1 and Algorithm 2 are well-suited for GPU implementation, where a single warp processes a single vector of size $M$. Fig. 3 illustrates the GPU implementation design for Algorithm 2, which can be divided into three stages: loading, searching, and selecting.
@@
Loading stage: In this stage, each vector is loaded from global memory into the corresponding shared memory, maintaining efficiency through coalescing memory access. A synchronization barrier is set at the end of this stage.
@@
Searching stage: In this stage, each vector is handled by a single warp, assuming the warp contains $w$ threads (Fig. 3 illustrates $w=4$, while in actual hardware environments $w=32$). The key point of the implementation is that the $max$, $min$, and $thres$ values at the beginning and in each iteration need to be synchronized across all threads within the warp. This can be achieved using a combination of classic tree-reduction and broadcast primitives. The first step is to obtain the $max$ and $min$ of the vector. The vector is divided into $\lceil M/w\rceil$ parts, with each thread responsible for extracting the maximum and minimum elements within its assigned part. Then, a tree-reduction using the shuffle primitive is performed in $\log_{2}{w}$ steps to obtain the maximum and minimum elements across the entire warp, and these values are broadcasted to all threads within the warp. The second step is to perform a binary search according to Algorithm 2 to obtain the selection threshold. In each iteration, the count of elements above the current threshold is accumulated and broadcasted using the same tree-reduction method. Then, each thread uses $cnt$ to simultaneously update the $max$, $min$, and $thres$ values. After a specified number of iterations, the final threshold is obtained.
Selecting stage: A single warp traverses the entire vector in one pass. The ballot primitive is used to identify the elements and their indices that meet the selection threshold, and these are dumped into the output buffer. The pop count primitive is then used to count the number of selected elements to ensure that only the top-k pairs are dumped.
@@
This design requires no data writes outside of registers, except for loading the vector and dumping the results. During the searching and selecting stages, warp-level primitives are utilized to achieve highly optimized inter-thread collaboration. Moreover, each warp operates independently in parallel, maintaining high overall efficiency. The implementation for Algorithm 1 follows the same workflow, with adjustments made to the termination condition of the loop in the searching stage. Additionally, the selecting stage may require a potential two-pass selection, which is accomplished by repeating the selection process with a different threshold.
@@ Appendix B A comprehensive analysis of the performance of RTop-K when applied to varying vector sizes
  • When $M$ is below 1280, RTop-K achieves more than a 5×5$\times$5 × speedup over PyTorch.
  • When $M$ is between 1280 and 3072, RTop-K achieves a 2×2$\times$2 × to 5×5$\times$5 × speedup over PyTorch.
  • When $M$ is between 3072 and 6144, RTop-K achieves a 1×1$\times$1 × to 2×2$\times$2 × speedup over PyTorch.

### v4

@@ 2.3 GPU Top-k Implementations
4:  $cnt\leftarrow 0$
5:  while $max-min>\epsilon$ do
6:     $thres\leftarrow\frac{min+max}{2}$
7:     $cnt\leftarrow\left|\{i\mid\mathbf{v}_{i}\geq thres\}\right|$
8:     if $cnt<k$ then
9:        $max\leftarrow thres$
10:     else if $cnt>k$ then
11:        $min\leftarrow thres$
12:     else
13:        break
14:     end if
@@
16:  if $cnt<k$ then
17:     $\mathbf{elems},\mathbf{indices}\leftarrow\{(\mathbf{v}_{i},i)\mid\mathbf{v}_{i}\geq thres\}$
18:     Append the first $k-cnt$ pairs of $\{(\mathbf{v}_{i},i)\mid min\leq\mathbf{v}_{i}<thres\}$ to $\mathbf{elems},\mathbf{indices}$
19:  else
20:     $\mathbf{elems},\mathbf{indices}\leftarrow$ the first $k$ pairs of $\{(\mathbf{v}_{i},i)\mid\mathbf{v}_{i}\geq thres\}$
21:  end if
22:  return  $\mathbf{elems},\mathbf{indices}$
@@ 3.1 Binary Search-based Top-k Selection Algorithm
The algorithm, as illustrated in Fig. 2, first retrieves the $min$ and $max$ values of the vector, and then uses several iterations of binary search to determine a threshold $thres$. The $min$, $max$, and $thres$ values are updated in each iteration, and the loop terminates when the number of elements filtered by the current threshold equals $k$. One corner case that must be considered is when $\mathbf{v}$ contains multiple equal or very close elements, and these elements happen to be near the ”borderline” during top-k selection. In this scenario, relying solely on $thres$ for selection makes it difficult or even impossible to extract exactly $k$ elements. It can be verified that in such cases, these borderline elements are always located between $min$ and $thres$. Therefore, a second filtering step using $min$ is sufficient to supplement the selection and ensure that exactly $k$ elements are chosen. The parameter $\epsilon$ represents the precision of the algorithm, which determines the maximum width of the borderline and imposes an upper bound on the number of iterations in the loop. When $\epsilon$ is set to 0 (which is effectively equivalent to the minimal distinguishable precision of the float type), Algorithm 1 becomes an exact top-$k$ selection algorithm.
@@ 3.2 GPU Implementation Design
Both Algorithm 1 and Algorithm 2 are well-suited for GPU implementation, where a single warp processes a single vector of size $M$. Fig. 3 illustrates the GPU implementation design, which can be divided into three stages: loading, searching, and selecting.
@@
Loading stage: In this stage, each vector is loaded from global memory into the corresponding shared memory, which is done by a warp looping through it. A synchronization barrier is set at the end of this stage.
@@
Searching stage: In this stage, each vector is handled by a single warp, assuming the warp contains $w$ threads (Fig. 3 illustrates $w=4$, while in actual hardware environments $w=32$). The key point of the implementation is that the $max$, $min$, and $thres$ values at the beginning and in each iteration need to be synchronized across all threads within the warp. This can be achieved using a combination of classic tree-reduction and broadcast primitives. The first step is to obtain the $max$ and $min$ of the vector. The vector is divided into $\lceil M/w\rceil$ parts, with each thread responsible for extracting the maximum and minimum elements within its assigned part. Then, a tree-reduction using the shuffle primitive is performed in $\log_{2}{w}$ steps to obtain the maximum and minimum elements across the entire warp, and these values are broadcasted to all threads within the warp. The second step is to perform a binary search to determine the selection threshold. In each iteration, the count of elements exceeding the current threshold is accumulated and broadcasted using the same tree-reduction method. Then, each thread uses $cnt$ to simultaneously update the $max$, $min$, and $thres$ values. Once the exit condition is met or the maximum number of iterations is reached, the final threshold is obtained.
Selecting Stage: A single warp performs a one-pass or two-pass selection over each vector. Let $cnt$ be the number of elements greater than or equal to the final threshold $thres$. The selection applies two conditions separately: selecting elements where $x\geq thres$, and, if needed, selecting additional elements where $thres>x\geq min$.
If $cnt\geq k$, the first condition is sufficient to produce $k$ elements, and the second condition is skipped. Otherwise, all elements satisfying the first condition are selected, and the remaining $k-cnt$ elements are supplemented using the second condition.
To implement this efficiently, we use the ballot primitive to merge and broadcast the selection masks across threads within a warp. The popcnt primitive is then used to compute the inclusive prefix sum of selected elements for each thread. Based on the prefix sum, each thread finalizes its selection decision: if the prefix sum exceeds $k$, its selection is disabled. The values and indices of the final selected elements are written to the output buffer in global memory.
@@
This design requires no data writes outside of registers, except for loading the vector and dumping the results. During the searching and selecting stages, warp-level primitives are utilized to achieve highly optimized inter-thread collaboration. Moreover, each warp operates independently in parallel, maintaining high overall efficiency.
@@ Appendix B A comprehensive analysis of the performance of RTop-K when applied to varying vector sizes
  • When $M$ is below 1280, RTop-K achieves a 4.9×4.9$\times$4.9 × to 12.5×12.5$\times$12.5 × speedup over PyTorch.
  • When $M$ is between 1280 and 3072, RTop-K achieves a 2.3×2.3$\times$2.3 × to 4.9×4.9$\times$4.9 × speedup over PyTorch.
  • When $M$ is between 3072 and 6144, RTop-K achieves a $\times1$ to 2.3×2.3$\times$2.3 × speedup over PyTorch.